OpenStudy (anonymous):

solve. log2 (x^2)-3=13

OpenStudy (anonymous):

+ or - 2?

OpenStudy (anonymous):

\[\log _{2}(x^2)-3=13 \]?

OpenStudy (anonymous):

or \[\log2(x^2)-3=13\]

OpenStudy (anonymous):

I read this as \[\log_{2} \] and came up with \[\pm\]256

OpenStudy (anonymous):

I've done this completely wrong!

OpenStudy (anonymous):

yes what sjsharks has is what i was looking for. ty